# Cross-Modal Enhancement and Benchmark for UAV-based Open-Vocabulary Object Detection

### v1

@@ Abstract
Open-Vocabulary Object Detection (OVD) has emerged as a pivotal technology for applications involving Unmanned Aerial Vehicles (UAVs). However, the prevailing large-scale datasets for OVD pre-training are predominantly composed of ground-level, natural images. This creates a significant domain gap, causing models trained on them to exhibit a substantial drop in performance on UAV imagery. To address this limitation, we first propose a refined UAV-Label engine. Then we construct and introduce UAVDE-2M(contains over 2,000,000 instances and 1800 categories) and UAVCAP-15k(contains over 15,000 images). Furthermore, we propose a novel Cross-Attention Gated Enhancement Fusion (CAGE) module and integrate it into the YOLO-World-v2 architecture. Finally, extensive experiments on the VisDrone and SIMD datasets verify the effectiveness of our proposed method for applications in UAV-based imagery and remote sensing.
@@ 1 Introduction
To effectively embed textual guidance into UAV visual representations, we propose the cross-attention gated enhancement(CAGE) fusion module. The module’s innovation lies in a dual-path architecture that leverages a cross-attention governed by an gate mechanism to ensure precise, spatially-aware semantic grounding, while a parallel FiLM layer provides global feature modulation, with both pathways integrated into the visual stream via a residual connection for robust enhancement. We summarize the main contributions as follows,
@@ 3.1 UAVDE-2M
Building upon the LAE-Label Engine, our UAV-Label Engine introduces a more sophisticated data handling pipeline. Its core innovation is the classification of data into a new category—Partially-annotated Datasets (FOP)—alongside the standard FOD and COD. This FOP category is specifically designed to address datasets with valuable but not immediately usable annotations, tackling common challenges such as: 1) frame redundancy in UAV video data, 2) incompatible annotation types (e.g., segmentation vs. rotated boxes), and 3) ambiguous class labels (e.g., ‘vehicle,’ ‘obstacle’).
@@ 4.1 Cross-Attention Gated Enhancement Module
To effectively integrate natural language guidance with visual features, we propose a novel fusion module, termed CAGE. It enriches spatial image features with semantic context derived from textual descriptions. The module employs a multi-stage architecture combining: (1) a cross-modal attention mechanism, (2) a spatial context refinement block with an adaptive gate, (3) a global FiLM-based conditioning layer, and (4) a residual connection. It operates on an image feature map $F_{img}$ and text embeddings $F_{text}$. The detail structure of CAGE is shown Figure 6.
The cornerstone of our module is a multi-head cross-attention mechanism that generates a text-aware context map. Visual features $F_{img}$ are linearly projected to form the query tensor $Q$. Concurrently, the text embeddings $F_{text}$ are transformed into key ($K$) and value ($V$) tensors, using two distinct linear projections. Each spatial location in the image feature map (a query) attends to all text tokens (keys) to compute attention scores. These scores are then used to compute a weighted sum of the text values ($V$), yielding a context vector for each spatial location. The attention output, which forms a context map $F_{ctx}$, is computed as:
$F_{ctx}=softmax\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
where $d_{k}$ is the dimension of each attention head.
The raw context map $F_{ctx}$ is first refined through a lightweight depthwise separable convolution block to enhance local spatial coherence. To prevent visual information from being overwhelmed, we introduce an gating mechanism. This gate, implemented as a small convolutional head, generates a spatial mask $M_{\text{occ}}$ from $F_{img}$ to adaptively modulate the intensity of the refined context map ($F^{\prime}_{ctx}\odot M_{occ}$). The gated context is then concatenated with the original image features $F_{img}$ and merged via a 1x1 convolution.
As a parallel path for global conditioning, we employ Feature-wise Linear Modulation (FiLM). Global average-pooled text features are passed through an MLP to predict channel-wise scale ($\gamma$) and bias ($\beta$) parameters. These parameters modulate the merged features:
$F_{out}=F_{merged}\odot(1+\gamma)+\beta$ (2)
Finally, a residual connection from the input $F_{img}$ is added to the modulated output to preserve the original visual information pathway, followed by a Batch Normalization layer to stabilize the final output.
@@
4.2 Module Embedding
To validate the efficacy and versatility of our proposed Cross-Attention Gated Enhancement (CAGE) module, we integrate it into the state-of-the-art real-time open-vocabulary detection framework, YOLO-World-v2. Our integration strategy is designed to be seamless, positioning the CAGE module as a extitdrop-in replacement for the original text-vision fusion blocks, thereby demonstrating its utility as a flexible, plug-and-play component for various OVD models.
Integration Strategy.
The core architecture of YOLO-World-v2 features a sophisticated neck structure, derived from the Path Aggregation Network (PAN), which is responsible for fusing multi-scale visual features. Within this neck, the T-CSPLayer (Text-guided Cross-Stage Partial Layer) serves as the primary component for injecting textual guidance into the visual processing pipeline. We identify these T-CSPLayers as the ideal integration points.
Our strategy involves replacing every instance of the T-CSPLayer in the YOLO-World-v2 neck with our CAGE module. This ensures that text-vision fusion is consistently handled by our proposed method across all feature scales (e.g., P3, P4, P5). The process is facilitated by a strict alignment of the input/output (I/O) dimensions:
  • extbfInput: For each CAGE module, the visual feature map $F_{img}$ is taken directly from the corresponding level of the PAN. The text embeddings $F_{text}$ are supplied by the model’s shared text encoder (e.g., CLIP text encoder).
  • extbfOutput: Our CAGE module is carefully configured to produce an output tensor that has the exact same channel count and spatial resolution as the T-CSPLayer it replaces. This ensures that the data flow to subsequent layers in the YOLO-World-v2 neck and head remains undisturbed, requiring no further architectural modifications.
Rationale and Advantages.
This substitution is motivated by the advanced capabilities of our module. While the T-CSPLayer provides effective fusion, our CAGE module offers a more nuanced mechanism. The multi-head cross-attention facilitates a finer-grained alignment between specific visual regions and semantic concepts in the text. Furthermore, the inclusion of the extitocclusion gate and the global FiLM conditioning path provides a more robust fusion that preserves critical visual details while adaptively leveraging linguistic guidance. By implementing this as a drop-in replacement, we minimize architectural disruption, allowing us to fairly evaluate the direct performance impact of our fusion design. This modularity also highlights the potential for the CAGE module to be easily adapted to other vision-language architectures that require sophisticated cross-modal feature fusion.
@@ 5.3 Real World Application
To validate the performance of our proposed model on the real-world UAV-based open-vocabulary object detection task, as well as its real-time operational capabilities on edge computing devices, we deployed our proposed model on an NVIDIA Jetson Orin NX 16G embedded computing device. This device was mounted on a P600 UAV, enabling the system to capture the video feed from the UAV’s gimbal camera, perform inference in real-time, and subsequently transmit the detection results back to a ground station via a video transmission module. The image og our experimental platform is shown in Figure 1. Table 5 presents a comparison of the inference speeds between our proposed model and the baseline models on the NVIDIA Jetson Orin NX 16G. The results are shown in Figure 7. Our proposed method outperforms YOLO-World in terms of performance, as highlighted in the second column, our method demonstrates superior performance over YOLO-World-v2, particularly in challenging scenes with dense objects and heavy occlusion(e.g. parking lots).

### v2

# Light-Weight Cross-Modal Enhancement Method with Benchmark Construction for UAV-based Open-Vocabulary Object Detection

@@ Abstract
Open-Vocabulary Object Detection (OVD) faces severe performance degradation when applied to UAV imagery due to the domain gap from ground-level datasets. To address this challenge, we propose a complete UAV-oriented solution that combines both dataset construction and model innovation. First, we design a refined UAV-Label Engine, which efficiently resolves annotation redundancy, inconsistency, and ambiguity, enabling the generation of large-scale UAV datasets. Based on this engine, we construct two new benchmarks: UAVDE-2M, with over 2.4M instances across 1,800+ categories, and UAVCAP-15K, providing rich image-text pairs for vision-language pretraining. Second, we introduce the Cross-Attention Gated Enhancement (CAGE) module, a lightweight dual-path fusion design that integrates cross-attention, adaptive gating, and global FiLM modulation for robust text–vision alignment. By embedding CAGE into the YOLO-World-v2 framework, our method achieves significant gains in both accuracy and efficiency, notably improving zero-shot detection on VisDrone by +5.3 mAP while reducing parameters and GFLOPs, and demonstrating strong cross-domain generalization on SIMD. Extensive experiments and real-world UAV deployment confirm the effectiveness and practicality of our proposed solution for UAV-based OVD.
@@ 1 Introduction
To effectively embed textual guidance into UAV visual representations, we propose the cross-attention gated enhancement(CAGE) fusion module. The module’s innovation lies in a dual-path architecture that leverages a cross-attention governed by an gate mechanism to ensure precise, spatially-aware semantic grounding, while a parallel FiLM layer provides global feature modulation, with both pathways integrated into the visual stream via a residual connection for robust enhancement.
Finally, to validate the practical efficacy of our model and dataset in real-world UAV applications, we deployed the trained model onto our custom UAV experimental platform and conducted a series of object detection experiments.We Figure 1 illustrates the UAV platform.
We summarize the main contributions as follows,
@@ 3.1 UAVDE-2M
Building upon the LAE-Label Engine, our UAV-Label Engine introduces a more sophisticated data handling pipeline. Its core innovation is the classification of data into a new category named Partially-annotated Datasets (FOP) which alongside the standard FOD and COD. This FOP category is specifically designed to address datasets with valuable but not immediately usable annotations, tackling common challenges such as: 1) frame redundancy in UAV video data, 2) incompatible annotation types (e.g., segmentation vs. rotated boxes), and 3) ambiguous class labels (e.g., ‘vehicle,’ ‘obstacle’).
@@ 4.1 Cross-Attention Gated Enhancement Module
To effectively integrate natural language guidance with visual features, we propose a novel fusion module, termed CAGE (Context-Aware Gated Ensemble). It enriches spatial image features with semantic context derived from textual descriptions. The module operates on an image feature map $F_{img}\in\mathbb{R}^{B\times H\times W\times C}$ and a sequence of text embeddings $F_{text}\in\mathbb{R}^{B\times N\times D}$, where $D$ denotes batch size, $H,W$ are spatial dimensions, $C$ denotes channel depth, $D$ donates the number of categories, and $D$ donates the depth of the text embeddings. The module employs a multi-stage architecture combining: (1) a cross-modal attention mechanism, (2) a spatial context refinement block with an adaptive gate, (3) a global FiLM-based conditioning layer, and (4) a residual connection. The detailed structure of CAGE is shown in Figure 6.
The fusion process initiates by generating a spatial-semantic context map via multi-head cross-attention, where a query tensor, $Q$, is derived from the image features. The features are first projected by a $1\times 1$ convolution, spatially flattened, and then normalized, a process which can be formulated as:
$Q={LayerNorm}({Flatten}({Conv}_{1\times 1}(F_{img})))$ (1)
where $P$ denotes the projection dimension. Currently, the key ($K$) and value ($V$) tensors are derived from the text features after layer normalization, followed by distinct linear transformations.
$F^{\prime}_{text}={LayerNorm}(F_{text})$ (2)
$K$ $=F^{\prime}_{text}W_{K}\in\mathbb{R}^{B\times L\times P}$ (3)
$V$ $=F^{\prime}_{text}W_{V}\in\mathbb{R}^{B\times L\times P}$
where $W_{K}$ and $W_{V}$ are the weight matrices of the linear projection layers. The attention scores are computed via scaled dot-product attention across $h$ heads, each with dimension $d_{k}=P/h$. For the $i$-th head, the operation is:
$head_{i}={softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{k}}}\right)V_{i}$ (4)
The outputs of all heads are concatenated and reshaped to form the context map ${ctx}_{map}\in\mathbb{R}^{B\times P\times H\times W}$.
In gated context refinement, the raw context map $ctx_{map}$ undergoes a refinement and gating process to enhance local features. The context map is first projected to an intermediate channel dimension and then refined by a block $\mathcal{F}_{dw}$ consisting of depthwise separable convolutions and GELU activations. An optional occlusion gate generates a per-pixel probability map $G\in[0,1]^{B\times 1\times H\times W}$ from the input image features, which determines the relevance of the textual context at each location, formulated as:
$G=\sigma(Conv_{1\times 1}(GELU(Conv_{3\times 3}(F_{img}))))$ (5)
where $\sigma$ is the Sigmoid function. The refined context is then modulated by this gate:
$ctx_{\text{ref}}=G\odot\mathcal{F}_{dw}(Conv_{1\times 1}(ctx_{map}))$ (6)
where $\odot$ denotes element-wise multiplication.
The final stage integrates the refined context with the original image features using both concatenation and global modulation. The gated context $ext{ctx}_{ext{ref}}$ is concatenated with the original image features $F_{img}$ along the channel dimension and subsequently fused via a convolutional block $\mathcal{F}_{merge}$ to produce a pre-modulated feature map:
$F_{pre-film}=\mathcal{F}_{merge}([F_{img};ctx_{ref}])$ (7)
For global semantic control, a FiLM layer is utilized. The text features $F_{text}$ are globally pooled to a sentence-level vector $f_{pool}$. This vector is passed through a multi-layer perceptron $\mathcal{F}_{film}$ to predict the channel-wise affine transformation parameters $\gamma$ and $\beta$:
$\gamma,\beta]=\mathcal{F}_{film}(f_{pool})\quad(\gamma,\beta\in\mathbb{R}^{B\times C_{out}\times 1\times 1})$ (8)
The pre-modulated features are then transformed as follows:
$F_{modulated}=(1+\gamma)\odot F_{pre-film}+\beta$ (9)
This operation allows the text command to scale and shift the visual feature activations dynamically.
To preserve the original visual information and ensure stable gradient flow, the module incorporates a parallel residual path, $F_{res}=\mathcal{F}_{res}(F_{img})$, where $\mathcal{F}_{res}$ is either a $1\times 1$ convolution or an identity map. The final output $F_{out}$ is the sum of the residual path and the batch-normalized modulated features:
$F_{out}=F_{res}+BatchNorm(F_{modulated})$ (10)
This design ensures that the fusion process enriches the visual features without catastrophically altering them, leading to a robust and effective fusion.
@@
4.2 Module Integration Strategy
To seamlessly integrate the proposed CAGE module, we replace the T-CSPLayer blocks in the YOLO-World-v2 neck with our design. Each CAGE module receives multi-scale visual features from the PAN structure and text embeddings from the shared text encoder. The output dimensions are carefully aligned with those of the replaced T-CSPLayer, ensuring that the downstream architecture remains unchanged. This plug-and-play design allows CAGE to serve as a direct substitute for existing text–vision fusion modules without introducing additional architectural overhead.
4.3 Rationale and Advantages
Compared with the original T-CSPLayer, CAGE provides a more fine-grained and robust text–vision alignment. Its multi-head cross-attention enhances local semantic grounding, the occlusion-aware gating mechanism adaptively filters irrelevant context, and the global FiLM pathway modulates features at the channel level. These enhancements collectively yield richer cross-modal representations while preserving essential visual details through residual connections. By adopting a drop-in replacement strategy, CAGE not only improves detection accuracy but also demonstrates strong generalizability, making it easily transferable to other vision–language frameworks.
@@ 5.3 Real World Application
To validate the performance of our proposed model on the real-world UAV-based open-vocabulary object detection task, as well as its real-time operational capabilities on edge computing devices, we deployed our proposed model on an NVIDIA Jetson Orin NX 16G embedded computing device. This device was mounted on a P600 UAV, enabling the system to capture the video feed from the UAV’s gimbal camera, perform inference in real-time, and subsequently transmit the detection results back to a ground station via a video transmission module. Table 5 presents a comparison of the inference speeds between our proposed model and the baseline models on the NVIDIA Jetson Orin NX 16G. The results are shown in Figure 7. Our proposed method outperforms YOLO-World in terms of performance, as highlighted in the second column, our method demonstrates superior performance over YOLO-World-v2, particularly in challenging scenes with dense objects and heavy occlusion(e.g. parking lots).